# DenoiseRep: Denoising Model for Representation Learning

### v2

@@ 3.1 Review Representation Learning
Representation learning plays a pivotal role in discriminative tasks, which is defined as "learning representations (or features) of the data that make it easier to extract useful information when building classifiers or other predictors" [4]. A common architecture of discriminative tasks consists of a vision backbone to extract discriminative features (e.g., ResNet [18], ViT [14]) and a task-specific head that operates on these features (e.g. MLP [24] for classification, RCNN [44] for object detection, FCN [long2015fully] for segmentation). It is evident that the vision backbone is central to representation learning. In this paper, we introduce a novel Denoising Model for Representation Learning (DenoiseRep), which integrates feature extraction and feature denoising within a single vision backbone. This approach aims to enhance the discriminative power of the features extracted.
@@ 4 Experiments
Detection Mask-RCNN [he2017mask] SwinV2-T COCO AP 42.80% 44.30%
Segmentation FCN [long2015fully] ResNet-50 ADE20K BIoU 28.70% 29.90%
@@ 4.2 Comparison with State-of-the-Art ReID Methods
CLIP-REID [li2023clip] ViT-base 75.80 89.70 90.50 95.40 83.10 90.80
@@ Appendix D Experiment on Image Detection Task
In this section, we aim to test the generalization ability of DenoiseRep in image detection tasks. We conduct experiments on the COCO [karpathy2015deep] dataset. The COCO (Common Objects in Context) dataset is a widely used dataset for large-scale image recognition, object detection, and image segmentation, particularly in computer vision tasks. It contains 80 types of objects, such as people, animals, daily necessities, etc., covering various common items in daily life. To verify that our method is model independent, we conduct experiments using different models including Mask-RCNN [he2017mask], Faster-RCNN [44], ATSS [zhang2020bridging], YOLO [redmon2018yolov3], DETR [carion2020end] and CenterNet [zhou2019objects], as well as diverse backbones. To evaluate the effectiveness of our method, we used standard metrics including AP, AP50, and AP75, which are commonly used metrics in object detection tasks to evaluate model performance, particularly for evaluating the effectiveness of bounding box detection. It is also an important part of the COCO dataset evaluation criteria, which can measure the detection ability of the model in multiple categories and scales.
@@ Appendix E Experiment on Image Segmentation Task
In this section, our objective is to assess the generalization capability of DenoiseRep in image segmentation tasks. The image segmentation task aims to divide an image into multiple regions in order to identify and understand objects or areas within them. The main challenges it faces include: complex and varied backgrounds that can easily interfere with segmentation results, objects obstruct each other, making segmentation difficult, objects have diverse shapes and may undergo deformations, etc. We conduct experiments on the ADE20K [zhou2017scene] dataset using the current mainstream image segmentation models. ADE20K is a widely used scene segmentation dataset, mainly used for image segmentation tasks. It contains approximately 20000 images and over 150 different object and region categories. We choose mIoU and B-IOU as evaluation metrics to comprehensively evaluate the performance of image segmentation models. MIoU is the average of IoUs for all categories, which can effectively reflect the segmentation ability of the model on different categories. It provides a quantitative model for the accuracy of handling complex scenes by calculating the degree of overlap between the predicted area and the real area. A higher mIoU value means that the model can better identify and segment the target object. B-IoU focuses on evaluating the accuracy of segmentation boundaries and is particularly suitable for object edge segmentation tasks. It provides sensitivity to boundary details by measuring the degree of overlap between predicted boundaries and real boundaries.
@@
Methods Backbones aAcc B-IoU mIoU Baseline +DenoiseRep Baseline +DenoiseRep Baseline +DenoiseRep FCN [long2015fully] ResNet-50 0.774 0.779 0.287 0.299 0.359 0.365 FCN ResNet-101 0.793 0.796 0.306 0.316 0.396 0.404 SegFormer [xie2021segformer] mit_b0 0.782 0.788 0.292 0.297 0.374 0.381 SegFormer mit_b1 0.812 0.816 0.341 0.348 0.422 0.425

### v4

@@ 3.1 Review Representation Learning
Representation learning plays a pivotal role in discriminative tasks, which is defined as "learning representations (or features) of the data that make it easier to extract useful information when building classifiers or other predictors" [4]. A common architecture of discriminative tasks consists of a vision backbone to extract discriminative features (e.g., ResNet [18], ViT [14]) and a task-specific head that operates on these features (e.g. MLP [26] for classification, RCNN [49] for object detection, FCN [40] for segmentation). It is evident that the vision backbone is central to representation learning. In this paper, we introduce a novel Denoising Model for Representation Learning (DenoiseRep), which integrates feature extraction and feature denoising within a single vision backbone. This approach aims to enhance the discriminative power of the features extracted.
@@ 4 Experiments
Detection Mask-RCNN [19] SwinV2-T COCO AP 42.80% 44.30%
Segmentation FCN [40] ResNet-50 ADE20K BIoU 28.70% 29.90%
@@ 4.2 Comparison with State-of-the-Art ReID Methods
CLIP-REID [32] ViT-base 75.80 89.70 90.50 95.40 83.10 90.80
@@ Appendix D Experiment on Image Detection Task
In this section, we aim to test the generalization ability of DenoiseRep in image detection tasks. We conduct experiments on the COCO [22] dataset. The COCO (Common Objects in Context) dataset is a widely used dataset for large-scale image recognition, object detection, and image segmentation, particularly in computer vision tasks. It contains 80 types of objects, such as people, animals, daily necessities, etc., covering various common items in daily life. To verify that our method is model independent, we conduct experiments using different models including Mask-RCNN [19], Faster-RCNN [49], ATSS [67], YOLO [48], DETR [6] and CenterNet [75], as well as diverse backbones. To evaluate the effectiveness of our method, we used standard metrics including AP, AP50, and AP75, which are commonly used metrics in object detection tasks to evaluate model performance, particularly for evaluating the effectiveness of bounding box detection. It is also an important part of the COCO dataset evaluation criteria, which can measure the detection ability of the model in multiple categories and scales.
@@ Appendix E Experiment on Image Segmentation Task
In this section, our objective is to assess the generalization capability of DenoiseRep in image segmentation tasks. The image segmentation task aims to divide an image into multiple regions in order to identify and understand objects or areas within them. The main challenges it faces include: complex and varied backgrounds that can easily interfere with segmentation results, objects obstruct each other, making segmentation difficult, objects have diverse shapes and may undergo deformations, etc. We conduct experiments on the ADE20K [73] dataset using the current mainstream image segmentation models. ADE20K is a widely used scene segmentation dataset, mainly used for image segmentation tasks. It contains approximately 20000 images and over 150 different object and region categories. We choose mIoU and B-IOU as evaluation metrics to comprehensively evaluate the performance of image segmentation models. MIoU is the average of IoUs for all categories, which can effectively reflect the segmentation ability of the model on different categories. It provides a quantitative model for the accuracy of handling complex scenes by calculating the degree of overlap between the predicted area and the real area. A higher mIoU value means that the model can better identify and segment the target object. B-IoU focuses on evaluating the accuracy of segmentation boundaries and is particularly suitable for object edge segmentation tasks. It provides sensitivity to boundary details by measuring the degree of overlap between predicted boundaries and real boundaries.
@@
Methods Backbones aAcc B-IoU mIoU Baseline +DenoiseRep Baseline +DenoiseRep Baseline +DenoiseRep FCN [40] ResNet-50 0.774 0.779 0.287 0.299 0.359 0.365 FCN ResNet-101 0.793 0.796 0.306 0.316 0.396 0.404 SegFormer [63] mit_b0 0.782 0.788 0.292 0.297 0.374 0.381 SegFormer mit_b1 0.812 0.816 0.341 0.348 0.422 0.425